Status Report 27 April 2002 (LMG02-03)

From LMG Chief Scientist, Jose Torres to NBP Chief Scientist, Peter Wiebe

 

Hi Pete,

 

It has been a while since the last update. We are currently at the shelfbreak vic survey station 12. We arrived here at 0100 24 April and have completed a couple of nice ADCP surveys, a couple of Moc 1's and a couple of Moc 10's with tandem HTI sampling. Late last night we started getting some high winds, 35-45, and we have not been able to deploy nets today. We are running an ADCP grid right now which will give us good coverage of our offshore station 1a and our shelf station, 1b. The winds have come down during the course of the day, but the sea has stayed pretty high. We are hoping to do some netting tomorrow and then we will most likely depart for our Avian pick up Saturday at midnight. Big change in faunal composition here. Lots of Euphausia triacantha, very few E. superba except for the huge amounts of larvae in the upper 25. None of the beaucoups catches of Pleuragramma either.

 

From Rothera we will head for station 4. You saw 10/10 pack at station 40? Was it thick enough to walk on? These and other questions are keeping us up nights.

 

Best to all, Jose